OpenStudy (anonymous):

One pipe fills a storage pool in 3 hours. A second pipe fills the same pool in 3 hours. When a third pipe is added and all three are used to fill the pool, it takes only 1 hour. Find how long it takes the third pipe to do the job alone.

OpenStudy (fellowroot):

Do you know the final answer?

OpenStudy (fellowroot):

Just so that I can check my work.

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }+\frac{ 1 }{ 3 }+\frac{ 1 }{x }=\frac{ 1 }{ 1 }\]

OpenStudy (anonymous):

multiply all by 3 and solve for x

OpenStudy (fellowroot):